I am guessing this is a quadratic equation. (x+y-z)(x-y+z)
How do I solve it? What is the name of this equation? What are the steps to solve it?
I am an exchange student in Japan and I don't understand completly when the teacher is teaching. Thank you very much.

Answer by stanbon(75887)

(x+y-z)(x-y+z)
Put it in the form (a+b)(a-b)
=(x+(y-z))(x-(y-z))
= (x^2 - (y-z)^2)
----------------
That might be the answer you want, or you may want to expand it to:
=x^2-(y^2-2yz+z^2)
=x^2-y^2-z^2+2yz
=====================
Enjoy your stay.
=====================
Cheers,
Stan H.